Help Save & Exi Complimentary angles 13 A projectile is launched at a 35 degree angle. What other launch angle will result in the same range? 0.88 points 25 References 65
Answer #1

Since range R =( u^2 Sin2@ )/g

Hence for the two equal range at same initial speed

Theta = 90-35 = 55degree .

Option first is correct .